Alan saves somewhere between $3 and $10 of his allowance each week, depending on the expenses that come up that week. If you write down the amount saved each week, what kind of sequence will you see?
Arithmetic. Geometric. Both. Or neither which one is correct

Answer by ahmadfara(1)

p=(n+1/100)
p=(14+1/100)
p=3/20